### v1

# Group-Aware Partial Model Merging for Children’s Automatic Speech Recognition

Abstract
Automatic Speech Recognition (ASR) for children remains challenging, primarily due to large acoustic variability and limited availability of training data. While supervised fine-tuning of adult pre-trained models has shown promise, it often fails to capture group-specific characteristics variations among children. To address this, we introduce GRoup-Aware PARtial model Merging (GRAPAM), a parameter-efficient approach that combines unsupervised clustering, partial fine-tuning, and model merging. Our approach adapts adult-pre-trained models to children by first grouping the children’s data based on acoustic similarity. Each group is used to partially fine-tune an adult pre-trained model, and the resulting models are merged at the parameter level. Experiments conducted on the MyST children’s speech corpus indicate that GRAPAM achieves a relative improvement of 6% of Word Error Rate (WER) , using the same amount of data, outperforming full fine-tuning while training fewer parameters. These results highlight the promise of model merging as a scalable and effective strategy for children’s ASR.
I Introduction
Recent progress in Automatic Speech Recognition (ASR) has been achieved through the introduction of large-scale Transformer-based architectures. These models have achieved state-of-the-art performance on numerous adult speech corpora, driven primarily by two key factors: the substantial increase in model complexity, reaching billions of parameters in some configurations [radford2023robust, hsu2021hubert, peng25c_interspeech], and the availability of large training datasets, with some exceeding one million hours of speech [koluguri2025granary]. Despite these advancements, using state-of-the-art ASR systems for children’s speech remains challenging, with performance consistently lagging behind that achieved for adult speech [sri_end2end, gelin2021endtoend, rolland2024exploring]. This disparity can be mainly attributed to the high acoustic variability characteristic of children’s speech, both between and within individual speakers. These variabilities primarily arise from the development of the child’s speech production system, leading to fluctuations in fundamental frequency, shifting of formant structures, and substantial variation in both spectral and temporal features of speech [Acoustic_change_children, reviewASRchildren]. In addition, children’s limited phonetic and linguistic maturity further impedes accurate speech recognition [language_children2], posing an additional challenge for existing ASR systems.
An additional challenge is the limited availability of large-scale and diverse children’s speech datasets, which significantly hinders the development of robust ASR models trained from scratch [sri_end2end, gelin2021endtoend]. In response, prior works have investigated various data augmentation strategies designed to expand and diversify existing corpora. These include techniques such as pitch normalisation [pitchnorm, yadav2019significance], Vocal Tract Length Normalisation (VTLN)[VTLN2, patel2024improving], multi-task learning [rolland2022multilingual], and the incorporation of synthetically generated speech data [wang2021towards, kadyan2021synthesis, rolland2024improved, zhao23c_interspeech].
A widely adopted approach for improving children’s ASR performances is Supervised Fine-Tuning (SFT) [sri_end2end, gelin2021endtoend, fan2022towards, fan2024benchmarking, jain2023adaptation]. In this approach, the parameters of a pre-trained ASR model, originally trained on a large adult speech corpus are adapted to children’s speech via additional supervised training. By leveraging the general linguistic and acoustic representations learned during pretraining, SFT facilitates the effective transfer of knowledge to the children’s speech domain. Importantly, prior studies have demonstrated that it is often unnecessary to adapt the entire model, as fine-tuning only a subset is often sufficient to achieve comparable performance improvements [rolland2024introduction]. One of the principal advantages of SFT lies in its ability to achieve substantial gains in recognition accuracy with relatively limited training data, making it particularly suitable for low-resource child speech settings.
Nonetheless, a significant challenge of SFT for children’s ASR lies in the substantial acoustic variation across different age groups [1255448, gerosa2009review]. As children’s speech undergoes developmental changes with age, applying a uniform fine-tuning strategy across all age ranges may fail to capture age-specific acoustic characteristics effectively. Recent findings support the use of age-specific ASR models, which consistently yield higher recognition accuracy [lu2022improving, lilles2025fine, hamalainen2014improving]. Howevever, as there is one model per group, these models typically require prior knowledge of the speaker’s age, which is not always available, and raise concerns surrounding the scalability and parameter efficiency of storing and managing multiple distinct model copies.
Recently, model merging has emerged as a lightweight way to combine specialised models trained on different domains without joint retraining, particularly in the domain of large language models (LLMs) [yang2024model, wortsman2022model, yu2024language, jin2022dataless]. A key insight from recent work on model merging for LLMs is that the resulting merged model often preserves the capabilities of the individual constituent models [yu2024language, jin2022dataless]. While parameter-level model merging remains a relatively novel area of research, particularly within LLMs and Vision Language Models (VLMs) [chen2025bring, biggs2024diffusion], emerging studies have begun to explore its potential applicability in ASR [ducorroy25_interspeech, tan2021novo], including promising developments on children’s speech [shankar2025selective].
In this work, we explore the potential of model merging as a parameter-efficient strategy to improve ASR for children, with a particular focus on combining models fine-tuned on distinct groups of child speech. The key contributions of this work are as follows:
  1. 1.
    We introduce Group-Aware Partial Model Merging (GRAPAM), a novel framework that integrates unsupervised clustering, partial fine-tuning, and model merging to improve children’s ASR.
  2. 2.
    We conduct a comprehensive analysis of clustering and fine-tuning strategies for effective model merging.
  3. 3.
    We demonstrate consistent improvements over conventional fine-tuning for children’s ASR.
  4. 4.
    We propose and evaluate heterogeneous and iterative merging variants to further enhance performance.
II Related work
II-A Model merging
In recent years, LLMs and VLMs have demonstrated improved capabilities in tasks such as zero-shot inference and general-purpose understanding [radford2019language, brown2020language]. Owing to their training on diverse data, these models exhibit significant flexibility and generalisation capabilities. Nonetheless, several challenges persist. These include the generation of harmful or biased content [navigli2023biases], substantial high computational demands [10756602] and financial costs associated with training [li2023flm], and underperformance on specialised tasks [anjum2025domain].
In this context, model merging has emerged as a promising research direction to overcome these limitations. It aims to combine multiple task-specific models into a single unified architecture that preserve the capabilities of each constituent model [wortsman2022model, matena2022merging, ilharco2022editing]. Unlike traditional multi-task learning, which relies on joint optimisation over multiple datasets, model merging operates directly at the parameter level, enabling the integration of pre-trained models without access to the original training data or additional retraining.
A variety of model merging strategies have been introduced in recent literature. Average Merging, or Linear Interpolation (LERP) [wortsman2022model], constructs a merged model by directly averaging the parameters of multiple fine-tuned models. Task Arithmetic [ilharco2022editing] computes task-specific vectors by subtracting a shared base model from each task-specific model; these vectors are then combined using predefined scaling factors to control the relative contribution of each task. Fisher Merging [matena2022merging] performs weighted parameters fusion, where the weights are derived from the Fisher Information Matrix, emphasising parameters that are more informative. RegMean [jin2022dataless] formulates the merging process as a linear regression problem with a closed-form solution. More recently, TIES-Merging [yadav2023ties] was introduced to mitigate parameter conflicts inherent in task arithmetic by trimming low-magnitude parameters, resolving sign inconsistencies, and separately merging parameters with consistent signs. Finally, DARE Merging [yu2024language] can be applied in conjunction with other methods, randomly dropping a proportion of parameters and rescaling the remaining weights prior to merging to improve robustness.
Beyond the language and vision domains, model merging has also been explored in the context of ASR [ducorroy25_interspeech, tan2021novo]. An early approach, Divide-and-Merge (DAM) [tan2021novo], trained multiple models on different subsets of the dataset and subsequently combined them using genetic algorithms and SGD-based fine-tuning to merge the resulting acoustic models. More recently, model merging has been successfully applied to Whisper-based models fine-tuned for dysarthric speech recognition, demonstrating improved generalisation in low-resource and long-form scenarios [ducorroy25_interspeech]. Finally, Selective Attention Merging [shankar2025selective] introduced a layer-wise approach that merges attention-layer via task vectors between models trained on adult and child speech, yielding notable performance gains for children’s ASR.
II-B Partial Fine-tuning
As models continue to increase in scale, their parameter counts have grown dramatically. Consequently, full SFT of such large models is often impractical, especially in low-resource settings. In many cases, fine-tuning the entire model leads to suboptimal performance, as the vast number of trainable parameters can easily result in overfitting when only limited adaptation data is available. To address this challenge, parameter-efficient fine-tuning (PEFT) methods have emerged as compelling alternatives. Techniques such as Adapters [houlsby2019parameter, fan22d_interspeech, 10447091] and Low-Rank Adaptation (LoRA) [hu2021loralowrankadaptationlarge, 10447004] introduce additional trainable parameters while keeping the original pre-trained model weights frozen, thus enabling efficient and scalable adaptation. Another related approach, known as Partial Fine-Tuning (PFT) [ye2023partialfinetuningsuccessorfinetuning, Shen_Liu_Qin_Savvides_Cheng_2021, rolland2024exploring], involves updating only a subset of the existing parameters, typically within selected layers or components such as attention mechanisms or feed-forward networks, while leaving the rest of the model unchanged.
III Group-Aware Partial Model Merging
Figure 1: Overview of the four stages of the GRAPAM pipeline. With $\mathcal{D}$ the full dataset and $\theta_{0}$ the pre-trained adult model parameters. The fire icon denotes stages involving ASR fine-tuning.
In this work, we introduce GRAPAM, a novel approach inspired by several complementary lines of research: the demonstrated benefits of age-dependent ASR models [lu2022improving, lilles2025fine, hamalainen2014improving], the emerging success of model merging in children’s ASR [shankar2025selective], the divide-and-merge paradigm introduced in DAM [tan2021novo] and recent advances in partial fine-tuning techniques [rolland2024introduction]. GRAPAM is designed to improve performance on children’s speech by addressing speaker variability in a parameter-efficient manner through model merging. The method comprises four core stages, as illustrated in Figure 1.
Let $\mathcal{D}$ denote the training dataset of size $N$, defined as $\mathcal{D}={(x_{i},y_{i})}_{i=1}^{N}$, where $x_{i}$ represents the input speech signal and $y_{i}$ denotes the corresponding transcription.
First, $\mathcal{D}$ is partitioned into similarity-based groups. Although age would be a natural criterion, in most children’s datasets, age information is not available. Moreover, chronological age may not accurately reflect developmental maturity. To address this, we adopt an unsupervised clustering strategy inspired by prior work [rolland2024exploring]. We extract utterance embeddings $E=\{e_{1},e_{2},\ldots,e_{N}\}$, where a single vector represent each utterance. In a second stage, we apply a clustering algorithm to divide these embeddings into K groups. A standard clustering algorithm such as k-means may be employed for this purpose:
$\min_{\{C_{1},\ldots,C_{K}\}}\sum_{i=1}^{K}\sum_{x\in C_{i}}\|x-\mu_{i}\|^{2},$ (1)
where $\mu_{i}$ is the centroid of cluster $C_{i}$. Each cluster $C_{k}$ corresponds to a subset $\mathcal{D}_{k}\subset\mathcal{D}$ of the training data.
In the third stage, we perform PFT on each group-specific dataset $\mathcal{D}_{k}$ as well as on the entire dataset $\mathcal{D}$. Given a pre-trained ASR model with parameters $\theta_{0}$, we selectively fine-tune the feed-forward (FFN) or attention (ATTN) submodules across all layers of the Transformer architecture, resulting in
$\theta_{k}=\text{PFT}(\theta_{0},\mathcal{D}_{k}),\quad k=1,\ldots,K$ (2)
and
$\theta_{all}=\text{PFT}(\theta_{0},\mathcal{D})$ (3)
This selective fine-tuning strategy not only ensures parameter efficiency but also facilitates effective adaptation in low-resource scenarios. In this work, we opt for PFT over alternative PEFT approaches such as LoRA or Adapters, as PFT directly modifies existing model parameters. This property significantly simplifies parameter-level merging, which is central to our approach. Consequently, we focus on PFT leaving the exploration of GRAPAM with other PEFT strategies for future work.
In the final stage, we apply LERP to merge the independently fine-tuned models into a single, unified model:
$\theta_{\text{GRAPAM}}=\alpha_{all}\theta_{all}+\sum_{k=1}^{K}\alpha_{k}\theta_{k}$ (4)
with constraints that
$\alpha_{all}+\sum_{i=1}^{K}\alpha_{i}=1,\quad\alpha_{all}\geq 0\text{ and }\alpha_{i}\geq 0$ (5)
In this work, we set the interpolation weights $\alpha_{i}$ uniformly across all selected fine-tuned models, thereby ensuring that each model contributes equally to the final merged model:
$\alpha_{i}$ $=\begin{cases}\frac{1}{|\mathcal{M}|},&\text{if }i\in\mathcal{M}\\ 0,&\text{otherwise}\end{cases}$
$\text{for all }i\in\{all,1,\ldots,K\}$ (6)
with $\mathcal{M}\subseteq\{\theta_{all},\theta_{1},\dots,\theta_{k}\}$ the set of selected fine-tuned models for merging. While we use a uniform weight merge strategy, we note that $\alpha_{i}$ could be further optimised based on validation performance in future work. Finally, the resulting merged model, with parameters denoted as $\theta_{\text{GRAPAM}}$, is used for inference on the entire children’s test set. This unified model is expected to more effectively capture speaker-group-specific characteristics, thereby improving recognition performance.
IV Experimental setup
IV-A Corpus
TABLE I: My Science Tutor Children Speech Corpus statistics
Training Validation Test
# of utterances 42790 6812 7257
# of speakers 566 80 92
# of hours 117 18 19
For our experiments, we used the My Science Tutor Corpus (MyST) dataset. MyST is one of the largest publicly accessible collections of English children’s speech, comprising approximately 400 hours. It encompasses dialogues between children and a virtual tutor across eight scientific domains, involving 1,372 students in grades three to five. The corpus is pre-partitioned, ensuring equitable representation of scientific domains and unique student occurrences within each partition. However, only 45% of utterances are transcribed at the word level. In this work, we excluded utterances shorter than one second, mainly containing silence, and those longer than 20 seconds, due to GPU constraints. After this filtering process, the dataset comprises 56,859 utterances from 738 speakers, totaling approximately 154 hours of speech. More detailed statistics of the different partitions are provided in Table I.
IV-B Implementation details
All experiments were conducted using the SpeechBrain toolkit [ravanelli2021speechbrain]. For the ASR model, we employed Whisper-medium [^1] https://huggingface.co/openai/whisper-medium.en [radford2023robust], which has previously demonstrated strong performance when fine-tuned on children’s speech [jain23_interspeech]. Whisper-medium is a Transformer-based encoder-decoder architecture comprising approximately 763.9M parameters, with 24 encoder layers and 24 decoder layers. The model was originally trained on a large-scale, multilingual dataset consisting of approximately 680k hours of transcribed speech.
In this work, we evaluate three types of utterance-level representations for clustering. First, we extract 192-dimensional speaker embeddings using a pre-trained ECAPA-TDNN model [^2] https://huggingface.co/speechbrain/spkrec-ecapa-voxceleb. Second, we compute a set of 100-dimensional low-level speech metrics using the Librosa library [mcfee2015librosa], including pitch, pause rate, speaking rate, spectral centroid, and signal-to-noise ratio; these features are subsequently reduced to 16 dimensions using Principal Component Analysis (PCA). Third, we introduce a one dimensional WER-based representation derived from zero-shot inference WERs using the Whisper model, clustering utterances based on their predicted transcription quality.
For all three representation types, we apply the $k$-means clustering algorithm from scikit-learn [pedregosa2011scikit], with a fixed number of clusters. We observed that, despite employing k-means, the cluster sizes remain balanced across all methods. Unless stated otherwise, we use $K=3$ clusters throughout our experiments. As for comparison, we also include random clustering, which assigns utterances to clusters uniformly at random.
For Partial Fine-Tuning (PFT), we experiment with three configurations: fine-tuning (i) all model parameters (763.9M), (ii) only the attention components (302.4M), and (iii) only the feed-forward network (402.9M) (FFN) components across all Transformer layers. All experiments are conducted on a single NVIDIA RTX A6000 GPU (48 GB), with a batch size of 16, a learning rate of $1\times 10^{-5}$, and one epoch of training. The models are optimised using the negative log-likelihood (NLL) loss function.
Finally, for model merging, we used LERP with uniform interpolation of the different models present in the configuration, ensuring that each group of children contributed equally to the final merged model.
V Results and discussion
To comprehensively assess the effectiveness of GRAPAM, we conduct a series of ablation studies aimed at disentangling the contributions of its core components. In particular, we evaluate the impact of different utterance clustering strategies, the effect of varying the number of groups, the number of selected merged models, the role of PFT, and alternative merging strategies including iterative model merging. These experiments are presented and analysed in detail in the remainder of this section.
It is important to note that this study does not explore the influence of model size, using self-supervised learning (SSL) models instead of Whisper, or the volume of training data. These factors have already been examined in prior work on model merging for children’s ASR [shankar2025selective].
V-A Group-aware model merging
TABLE II: Results of different combinations for GRAPAM evaluated in WER (%). Values highlighted in grey indicate improvement relative to full-model fine-tuning on the entire dataset. The best-performing combinations are shown in bold.
$\mathcal{M}$ Clustering method
$\theta_{all}$ $\theta_{1}$ $\theta_{2}$ $\theta_{3}$ Spk-emb Random ZS WER LSM
14.05
x 9.95
x 10.34 10.11 10.25 9.73
x 9.79 9.69 10.32 10.61
x 10.36 10.63 10.18 10.03
x x 9.48 9.91 10.52 10.12
x x 9.77 10.34 10.09 9.45
x x 9.41 9.74 9.88 9.84
x x x 9.68 9.68 10.01 9.41
x x 9.72 9.98 10.32 9.79
x x 9.41 9.67 9.86 9.83
x x 9.76 9.77 9.96 9.70
x x x 9.36 9.64 9.98 9.97
x x x 9.68 9.98 9.72 9.68
x x x x 9.65 9.65 9.93 9.63
Average 9.59 9.84 10.03 9.74
To evaluate the effectiveness of our proposed approach, we first examine whether clustering the training dataset followed by model merging improves performance in children’s ASR, and which utterance embedding strategy yield the best results. As shown in Table II, the baseline system achieves a Word Error Rate (WER) of 14.05%. Fine-tuning the full model on the entire dataset significantly reduces the WER to 9.95%. This result confirms that SFT effectively mitigates the adult-speech bias inherent in the original pre-trained Whisper model.
Next, we partition $\mathcal{D}$ by group using several clustering strategies and apply model merging. For each configuration of $\mathcal{M}$, the different selected fine-tuned models used for merging are listed in the leftmost columns of Table II. Among all evaluated methods, grouping using speaker-embedding (Spk-emb) achieved the overall strongest performance, with an average WER of 9.59% and the lowest WER configuration with a score of 9.36%. The Low-level Speech Metrics (LSM) embeddings clustering also performs well but with less consistent results, reaching an average WER of 9.74% and a best configuration at 9.41%. Random clustering achieves 9.84% on average and 9.41% at best. Meanwhile, the Zero-Shot (ZS) WER clustering performs the least effectively with an average of 10.03% WER and best at 9.86%, showing no improvement over the baseline.
In the configurations where all the fine-tuned models trained on all the different data clusters are merged together (final row of Table II), or $\mathcal{M}=\{\theta_{all},\theta_{1},\theta_{2},\theta_{3}\}$, Spk-emb and Random embedding yield slightly higher WERs (9.65%) compared with LSM (9.63%). Nevertheless, all of them outperform the fine-tuning baseline, indicating that model merging remains an effective strategy.
Overall, these findings demonstrate that model merging enhances system performance relative to fine-tuning on the entire dataset and in particular, group-aware with speaker-embedding achieves the strongest and most consistent improvements.
V-B Influence of the number of groups
TABLE III: Influence of the number of clusters combination of all fine-tuned model ($\mathcal{M}=\{\theta_{all},\theta_{1},\dots,\theta_{k}\}$) and the best combination of $\mathcal{M}$ and results in WER (%).
Number of clusters Combination of all Best combination
1 9.95 9.95
2 9.42 9.41
3 9.65 9.36
4 9.64 9.33
We next evaluate model merging with varying numbers of clusters. Table III presents both the best-performing configurations of $\mathcal{M}$ and the results obtained by merging all fine-tuned group-based models ($\mathcal{M}=\{\theta_{all},\theta_{1},\theta_{2},\theta_{3}\}$) using the Spk-emb clustering method. The single-cluster setting corresponds to training on the entire dataset. For the best configurations, performance improves consistently as the number of clusters increases, with the lowest WER achieved using four clusters with a score of 9.33%. When all group-based models are merged, the combined system also delivers strong results, remaining competitive with the best configuration performance, achieving 9.42% with two clusters. These findings suggest that increasing cluster granularity allows the model to more effectively capture speaker-specific acoustic characteristics, leading to improved ASR performance when selected the right models to merge.
V-C Partial model merging
TABLE IV: WER (%) results of GRAPAM with partial fine-tuning. Grey indicates performance better than both full and partial fine-tuning on $\mathcal{D}$; Diagonal lines pattern indicates improvement over full fine-tuning but not over partial fine-tuning; The best-performing combination is shown in bold.
$\mathcal{M}$ Fine-tuning method
$\theta_{all}$ $\theta_{1}$ $\theta_{2}$ $\theta_{3}$ Full SFT PFT FFN PFT ATTN
x 9.95
Partial 9.48 9.46
x 10.34 10.03 10.00
x 9.79 9.68 9.68
x 10.36 9.58 9.71
x x 9.48 9.52 9.61
x x 9.77 9.69 9.68
x x 9.41 9.44 9.56
x x x 9.68 9.45 9.60
x x 9.72 9.66 9.66
x x 9.41 9.32 9.62
x x 9.76 9.66 9.66
x x x 9.36 9.33 9.68
x x x 9.68 9.32 9.66
x x x x 9.65 9.31 9.38
Partial x x x 9.36 9.51
Average 9.59 9.47 9.60
Table IV reports the WER performance of GRAPAM under three fine-tuning strategies: full model fine-tuning (Full SFT), partial fine-tuning of feed-forward modules (PFT FFN), and partial fine-tuning of attention modules (PFT ATTN). When applied to the entire dataset $\mathcal{D}$, full fine-tuning yields a WER of 9.95%. In contrast, partial fine-tuning achieves superior performance, with WERs of 9.48% (FFN) and 9.46% (ATTN), despite involving fewer trainable parameters.
Applying group-aware model merging further improves performance across all configurations, achieving the lowest WERs of 9.31% with PFT FFN and 9.38% with PFT ATTN. On average, PFT FFN delivers the strongest results (9.47%), outperforming both Full SFT (9.59%) and PFT ATTN (9.60%). Furthermore, we examine whether merging fully fine-tuned models or partially fine-tuned models, both trained on the entire data set, yields greater benefits within the GRAPAM framework. As shown in the final two rows of Table IV, full fine-tuning yields slightly better performance, achieving a WER of 9.31%, compared to 9.36% for PFT–FFN and 9.51% for PFT–ATTN. Interestingly, the best-performing PFT configurations correspond to merging all fine-tuned models, i.e., $\mathcal{M}=\{\theta_{\text{all}},\theta_{1},\theta_{2},\theta_{3}\}$.
V-D Integration of Clustering and Fine-Tuning Strategies
In this section, we explore the integration of the previously discussed clustering methods and fine-tuning strategies, assessing whether performance can be enhanced by merging models across different strategies or by combining them sequentially across training stages.
V-D1 Heterogeneous merging
Table V presents the WER results of GRAPAM for all combination of PFT and utterance embedding when merging all models from all clusters, where $\mathcal{M}=\{\theta_{all},\theta_{1},\theta_{2},\theta_{3}\}$. The table also includes the heterogeneous merging scenarios (MERGE ALL line and column), where models fine-tuned using different embedding types and fine-tuning strategies are combined together. The results indicate that PFT–FFN consistently yields the strongest performance with a score of 9.36% WER. Merging over utterance embedding strategy also performs competitively, recording a WER of 9.38% for both speaker embeddings and 9.40% for LSM and 9.39% for random clustering. Importantly, merging across the different utterance embeddings or fine-tuning strategies does not significantly improves or degrades performance, suggesting that the merging process effectively preserves model capabilities.
Notably, the MERGE ALL configuration, which merges all models across clustering types and fine-tuning strategies, achieves a WER of 9.32%, matching the best individual configurations and demonstrating the robustness of the GRAPAM framework under heterogeneous integration.
TABLE V: WER (%) results of different combination of PFT and utterance embedding as well as the heterogeneous merging of them. The best-performing combination is shown in bold.
Full SFT PFT FFN PFT Attn MERGE ALL
Spk embedding 9.65 9.31 9.38 9.38
LSM 9.63 9.32 9.45 9.40
Random 9.65 9.64 9.38 9.39
MERGE ALL 9.60 9.36 9.51 9.32
V-D2 Iterative Group-Aware Partial Merging
TABLE VI: WER (%) for different embedding methods (Spk embedding, LSM, Random) across PFT (FFN, ATTN, FULL) and iterations. Selected models as a source for the next turn are presented in bold.
PFT Utterance embedding TURN 1 TURN 2 TURN 3
Baseline - 9.48 10.14 10.28
FFN Spk embedding 9.31 9.30 9.35
LSM 9.32 9.29 9.34
Random 9.64 9.32 9.63
ATTN Spk embedding 9.38 9.32 9.36
LSM 9.45 9.30 9.35
Random 9.38 9.33 9.35
FULL Spk embedding 9.65 9.32 9.69
LSM 9.63 9.28 9.33
Random 9.65 9.88 9.92
Table VI presents the WER results obtained from successive iterations of GRAPAM, where, at each step, the best-performing configuration from the previous iteration is used as the pre-trained model for subsequent merging. In this experiment we used $\mathcal{M}=\{\theta_{all},\theta_{1},\theta_{2},\theta_{3}\}$.
First, we observe that when the full model is trained for additional epochs on the entire dataset, signs of overfitting emerge, as indicated by an increase in WER from 9.48% to 10.28%.
In contrast, the Iterative GRAPAM results demonstrate that initialising with speaker-embedding clustering and partial fine-tuning of the FFN yields the lowest WER of 9.28%, achieved in iteration 2 using LSM-based clustering and full fine-tuning. We hypothesise that this improvement arises from the model’s exposure to heterogeneous information across different cluster types at each iteration, enabling the progressive development of richer and more generalisable representations.
However, beyond iteration 2, no further improvements are observed. This plateau suggests that the model has saturated the informative capacity of the available cluster partitions and may be beginning to overfit.
These findings confirm that iterative group-aware partial model merging effectively enhances ASR performance.
VI Conclusion and future work
In this work, we presented GRAPAM, a parameter-efficient framework for adapting adult-pretrained ASR models to children’s speech through group-aware partial model merging. By combining unsupervised clustering, partial fine-tuning, and parameter-level interpolation, GRAPAM achieves a WER reduction from 9.95% to 9.31% on the MyST corpus. This new fine-tuning strategy for children’s ASR achieves comparable or better performance without needing multi-task training or additional data.
Future work can extend GRAPAM in several directions. First, adaptive weighting, where interpolation coefficients are optimised on held-out validation data rather than set uniformly is a promising direction. Second, more advanced clustering techniques can be considered, potentially incorporating self-supervised speech representations to better capture age and speaker-related variability. Finally, GRAPAM could be extended to other domains like pathological or multilingual speech recognition.

### v2

@@
-Group-Aware Partial Model Merging for Children’s Automatic Speech Recognition
+Rolland Abad
 Abstract
-Automatic Speech Recognition (ASR) for children remains challenging, primarily due to large acoustic variability and limited availability of training data.
-While supervised fine-tuning of adult pre-trained models has shown promise, it often fails to capture group-specific characteristics variations among children.
-To address this, we introduce GRoup-Aware PARtial model Merging (GRAPAM), a parameter-efficient approach that combines unsupervised clustering, partial fine-tuning, and model merging.
-Our approach adapts adult-pre-trained models to children by first grouping the children’s data based on acoustic similarity.
+While supervised fine-tuning of adult pre-trained models for children's ASR has shown promise, it often fails to capture group-specific characteristics and variations among children.
+To address this, we introduce GRoup-Aware PARtial model Merging, a parameter-efficient approach that combines unsupervised clustering, partial fine-tuning, and model merging.
+Our approach adapts adult-pre-trained models to children by first grouping the children's data based on acoustic similarity.
 Each group is used to partially fine-tune an adult pre-trained model, and the resulting models are merged at the parameter level.
-Experiments conducted on the MyST children’s speech corpus indicate that GRAPAM achieves a relative improvement of 6% of Word Error Rate (WER) , using the same amount of data, outperforming full fine-tuning while training fewer parameters.
-These results highlight the promise of model merging as a scalable and effective strategy for children’s ASR.
-I Introduction
-Recent progress in Automatic Speech Recognition (ASR) has been achieved through the introduction of large-scale Transformer-based architectures.
-These models have achieved state-of-the-art performance on numerous adult speech corpora, driven primarily by two key factors: the substantial increase in model complexity, reaching billions of parameters in some configurations [radford2023robust, hsu2021hubert, peng25c_interspeech], and the availability of large training datasets, with some exceeding one million hours of speech [koluguri2025granary].
-Despite these advancements, using state-of-the-art ASR systems for children’s speech remains challenging, with performance consistently lagging behind that achieved for adult speech [sri_end2end, gelin2021endtoend, rolland2024exploring].
-This disparity can be mainly attributed to the high acoustic variability characteristic of children’s speech, both between and within individual speakers.
-These variabilities primarily arise from the development of the child’s speech production system, leading to fluctuations in fundamental frequency, shifting of formant structures, and substantial variation in both spectral and temporal features of speech [Acoustic_change_children, reviewASRchildren].
-In addition, children’s limited phonetic and linguistic maturity further impedes accurate speech recognition [language_children2], posing an additional challenge for existing ASR systems.
-An additional challenge is the limited availability of large-scale and diverse children’s speech datasets, which significantly hinders the development of robust ASR models trained from scratch [sri_end2end, gelin2021endtoend].
-In response, prior works have investigated various data augmentation strategies designed to expand and diversify existing corpora.
-These include techniques such as pitch normalisation [pitchnorm, yadav2019significance], Vocal Tract Length Normalisation (VTLN)[VTLN2, patel2024improving], multi-task learning [rolland2022multilingual], and the incorporation of synthetically generated speech data [wang2021towards, kadyan2021synthesis, rolland2024improved, zhao23c_interspeech].
-A widely adopted approach for improving children’s ASR performances is Supervised Fine-Tuning (SFT) [sri_end2end, gelin2021endtoend, fan2022towards, fan2024benchmarking, jain2023adaptation].
-In this approach, the parameters of a pre-trained ASR model, originally trained on a large adult speech corpus are adapted to children’s speech via additional supervised training.
-By leveraging the general linguistic and acoustic representations learned during pretraining, SFT facilitates the effective transfer of knowledge to the children’s speech domain.
-Importantly, prior studies have demonstrated that it is often unnecessary to adapt the entire model, as fine-tuning only a subset is often sufficient to achieve comparable performance improvements [rolland2024introduction].
-One of the principal advantages of SFT lies in its ability to achieve substantial gains in recognition accuracy with relatively limited training data, making it particularly suitable for low-resource child speech settings.
-Nonetheless, a significant challenge of SFT for children’s ASR lies in the substantial acoustic variation across different age groups [1255448, gerosa2009review].
+Experiments conducted on the MyST children's speech corpus indicate that GRAPAM achieves a relative WER improvement of 6%, using the same amount of data, outperforming full fine-tuning while training fewer parameters.
+keywords:
+Children's speech, Children's ASR, Model merging, ASR, Partial fine-tuning
+1 Introduction
+Recent advances in Automatic Speech Recognition (ASR) have been made possible by the use of large-scale models, with state-of-the-art results on adult speech corpora [radford2023robust, hsu2021hubert, peng25c_interspeech].
+However, performance on children’s speech remains consistently inferior to that on adult speech [sri_end2end, gelin2021endtoend, rolland2024exploring], primarily due to the pronounced acoustic variability of children’s productions both inter and intra-speaker, arising from ongoing physiological and articulatory development that affects fundamental frequency, formant patterns, and broader spectral–temporal characteristics [Acoustic_change_children, reviewASRchildren].
+This difficulty is exacerbated by children’s limited phonetic and linguistic maturity [language_children2].
+In addition, progress is further constrained by the scarcity of large, diverse children’s speech corpora, which limits the feasibility of training robust models from scratch [sri_end2end, gelin2021endtoend].
+Consequently, existing work has explored augmentation and adaptation strategies such as pitch normalisation [pitchnorm, yadav2019significance], Vocal Tract Length Normalisation (VTLN) [VTLN2, patel2024improving], multi-task learning [rolland2022multilingual], and the use of synthetic speech [wang2021towards, kadyan2021synthesis, rolland2024improved, zhao23c_interspeech].
+Among these, supervised fine-tuning (SFT) arise as a widely adopted strategy [sri_end2end, gelin2021endtoend, fan2022towards, fan2024benchmarking, jain2023adaptation], adapting an adult-pretrained ASR model to children’s speech via additional training and thereby transferring general acoustic–linguistic representations to children's specifics.
+Importantly, it was observed that comparable gains can often be obtained by updating only a subset of parameters rather than the full network [rolland2024introduction], making SFT particularly effective in low-resource child-speech settings where substantial accuracy improvements are achievable with limited data.
+Nonetheless, a significant challenge of SFT for children's ASR lies in the substantial acoustic variation across different age groups [1255448, gerosa2009review].
 As children’s speech undergoes developmental changes with age, applying a uniform fine-tuning strategy across all age ranges may fail to capture age-specific acoustic characteristics effectively.
 Recent findings support the use of age-specific ASR models, which consistently yield higher recognition accuracy [lu2022improving, lilles2025fine, hamalainen2014improving].
-Howevever, as there is one model per group, these models typically require prior knowledge of the speaker’s age, which is not always available, and raise concerns surrounding the scalability and parameter efficiency of storing and managing multiple distinct model copies.
-Recently, model merging has emerged as a lightweight way to combine specialised models trained on different domains without joint retraining, particularly in the domain of large language models (LLMs) [yang2024model, wortsman2022model, yu2024language, jin2022dataless].
+However, group-specific fine-tuning produce many checkpoints raising concerns surrounding the scalability and parameter efficiency of storing and managing multiple distinct model copies.
+Additionally, these models typically require prior knowledge of the speaker’s age, which is not always available, making this group-specific training unavailable therefore losing its associated benefits.
+Model merging has emerged as a alternative way to combine specialised models trained on different domains without joint retraining, particularly in the domain of large language models (LLMs) [yang2024model, wortsman2022model, yu2024language, jin2022dataless].
 A key insight from recent work on model merging for LLMs is that the resulting merged model often preserves the capabilities of the individual constituent models [yu2024language, jin2022dataless].
@@
   3. 3.
-    We demonstrate consistent improvements over conventional fine-tuning for children’s ASR.
+    We demonstrate consistent improvements over conventional fine-tuning for children's ASR.
   4. 4.
     We propose and evaluate heterogeneous and iterative merging variants to further enhance performance.
-II Related work
-II-A Model merging
-In recent years, LLMs and VLMs have demonstrated improved capabilities in tasks such as zero-shot inference and general-purpose understanding [radford2019language, brown2020language].
-Owing to their training on diverse data, these models exhibit significant flexibility and generalisation capabilities.
-Nonetheless, several challenges persist.
-These include the generation of harmful or biased content [navigli2023biases], substantial high computational demands [10756602] and financial costs associated with training [li2023flm], and underperformance on specialised tasks [anjum2025domain].
-In this context, model merging has emerged as a promising research direction to overcome these limitations.
-It aims to combine multiple task-specific models into a single unified architecture that preserve the capabilities of each constituent model [wortsman2022model, matena2022merging, ilharco2022editing].
+2 Related work
+2.1 Model merging
+Model merging has emerged as a promising research direction to combine multiple task-specific models into a single unified architecture that preserve the capabilities of each constituent model [wortsman2022model, matena2022merging, ilharco2022editing].
 Unlike traditional multi-task learning, which relies on joint optimisation over multiple datasets, model merging operates directly at the parameter level, enabling the integration of pre-trained models without access to the original training data or additional retraining.
-A variety of model merging strategies have been introduced in recent literature.
-Average Merging, or Linear Interpolation (LERP) [wortsman2022model], constructs a merged model by directly averaging the parameters of multiple fine-tuned models.
-Task Arithmetic [ilharco2022editing] computes task-specific vectors by subtracting a shared base model from each task-specific model; these vectors are then combined using predefined scaling factors to control the relative contribution of each task.
-Fisher Merging [matena2022merging] performs weighted parameters fusion, where the weights are derived from the Fisher Information Matrix, emphasising parameters that are more informative.
-RegMean [jin2022dataless] formulates the merging process as a linear regression problem with a closed-form solution.
-More recently, TIES-Merging [yadav2023ties] was introduced to mitigate parameter conflicts inherent in task arithmetic by trimming low-magnitude parameters, resolving sign inconsistencies, and separately merging parameters with consistent signs.
-Finally, DARE Merging [yu2024language] can be applied in conjunction with other methods, randomly dropping a proportion of parameters and rescaling the remaining weights prior to merging to improve robustness.
-Beyond the language and vision domains, model merging has also been explored in the context of ASR [ducorroy25_interspeech, tan2021novo].
-An early approach, Divide-and-Merge (DAM) [tan2021novo], trained multiple models on different subsets of the dataset and subsequently combined them using genetic algorithms and SGD-based fine-tuning to merge the resulting acoustic models.
-More recently, model merging has been successfully applied to Whisper-based models fine-tuned for dysarthric speech recognition, demonstrating improved generalisation in low-resource and long-form scenarios [ducorroy25_interspeech].
-Finally, Selective Attention Merging [shankar2025selective] introduced a layer-wise approach that merges attention-layer via task vectors between models trained on adult and child speech, yielding notable performance gains for children’s ASR.
-II-B Partial Fine-tuning
-As models continue to increase in scale, their parameter counts have grown dramatically.
-Consequently, full SFT of such large models is often impractical, especially in low-resource settings.
-In many cases, fine-tuning the entire model leads to suboptimal performance, as the vast number of trainable parameters can easily result in overfitting when only limited adaptation data is available.
-To address this challenge, parameter-efficient fine-tuning (PEFT) methods have emerged as compelling alternatives.
-Techniques such as Adapters [houlsby2019parameter, fan22d_interspeech, 10447091] and Low-Rank Adaptation (LoRA) [hu2021loralowrankadaptationlarge, 10447004] introduce additional trainable parameters while keeping the original pre-trained model weights frozen, thus enabling efficient and scalable adaptation.
-Another related approach, known as Partial Fine-Tuning (PFT) [ye2023partialfinetuningsuccessorfinetuning, Shen_Liu_Qin_Savvides_Cheng_2021, rolland2024exploring], involves updating only a subset of the existing parameters, typically within selected layers or components such as attention mechanisms or feed-forward networks, while leaving the rest of the model unchanged.
-III Group-Aware Partial Model Merging
+Recent work has proposed diverse model merging strategies.
+Average Merging, or Linear Interpolation (LERP) averages parameters across fine-tuned models [wortsman2022model], while Task Arithmetic combines task vectors formed by subtracting a shared base model [ilharco2022editing].
+Fisher Merging uses Fisher-information weights to emphasise informative parameters [matena2022merging], and RegMean casts merging as closed-form linear regression [jin2022dataless].
+More recent methods address interference and robustness: TIES trims small-magnitude updates and resolves sign conflicts before merging [yadav2023ties], and DARE randomly drops and rescales parameters as a pre-merge regulariser [yu2024language].
+Model merging has also been explored for ASR [ducorroy25_interspeech, tan2021novo].
+Divide-and-merge (DAM) trains models on data partitions and combines them via genetic search and SGD fine-tuning [tan2021novo].
+Recent studies apply merging to Whisper checkpoints for dysarthric speech, improving generalisation in low-resource and long-form settings [ducorroy25_interspeech], and propose Selective Attention Merging, a layer-wise approach that merges attention layers via task vectors for adult–child transfer, improving children’s ASR [shankar2025selective].
+2.2 Partial Fine-tuning
+As model scale increases, full SFT becomes increasingly costly and, in low-resource regimes, can degrade performance due to overfitting from updating hundreds of millions of parameters.
+Parameter-efficient fine-tuning (PEFT) addresses this by adapting models with limited trainable parameters: adapter-based methods [houlsby2019parameter, fan22d_interspeech, 10447091] and LoRA [hu2021loralowrankadaptationlarge, 10447004] add small trainable modules while keeping the backbone frozen.
+A closely related strategy, partial fine-tuning (PFT) [ye2023partialfinetuningsuccessorfinetuning, Shen_Liu_Qin_Savvides_Cheng_2021, rolland2024exploring], instead updates only selected subsets of the original parameters (e.g., attention or feed-forward blocks), leaving the remainder unchanged.
+3 Group-Aware Partial Model Merging
 Figure 1: Overview of the four stages of the GRAPAM pipeline.
 With $\mathcal{D}$ the full dataset and $\theta_{0}$ the pre-trained adult model parameters.
-The fire icon denotes stages involving ASR fine-tuning.
-In this work, we introduce GRAPAM, a novel approach inspired by several complementary lines of research: the demonstrated benefits of age-dependent ASR models [lu2022improving, lilles2025fine, hamalainen2014improving], the emerging success of model merging in children’s ASR [shankar2025selective], the divide-and-merge paradigm introduced in DAM [tan2021novo] and recent advances in partial fine-tuning techniques [rolland2024introduction].
-GRAPAM is designed to improve performance on children’s speech by addressing speaker variability in a parameter-efficient manner through model merging.
-The method comprises four core stages, as illustrated in Figure 1.
+The fire icon denotes stages with ASR fine-tuning.
+We introduce GRAPAM, a novel method that combines insights from several complementary research directions: prior evidence for the effectiveness of age-dependent ASR models [lu2022improving, lilles2025fine, hamalainen2014improving], recent progress on model merging for children’s ASR [shankar2025selective], the divide-and-merge paradigm of DAM [tan2021novo], and advances in partial fine-tuning for parameter-efficient adaptation [rolland2024introduction].
+GRAPAM targets children’s speech by mitigating speaker related variability using parameter-efficient fine-tuning followed by model merging.
+The approach consists of four stages, illustrated in Figure 1.
 Let $\mathcal{D}$ denote the training dataset of size $N$, defined as $\mathcal{D}={(x_{i},y_{i})}_{i=1}^{N}$, where $x_{i}$ represents the input speech signal and $y_{i}$ denotes the corresponding transcription.
 First, $\mathcal{D}$ is partitioned into similarity-based groups.
-Although age would be a natural criterion, in most children’s datasets, age information is not available.
+Although age would be a natural criterion, in most children's datasets, age information is not available.
 Moreover, chronological age may not accurately reflect developmental maturity.
@@
 A standard clustering algorithm such as k-means may be employed for this purpose:
-$\min_{\{C_{1},\ldots,C_{K}\}}\sum_{i=1}^{K}\sum_{x\in C_{i}}\|x-\mu_{i}\|^{2},$ (1)
+$\min_{\{C_{1},\ldots,C_{K}\}}\sum_{i=1}^{K}\sum_{x\in C_{i}}\lVert x-\mu_{i}\lVert^{2}$ (1)
 where $\mu_{i}$ is the centroid of cluster $C_{i}$. Each cluster $C_{k}$ corresponds to a subset $\mathcal{D}_{k}\subset\mathcal{D}$ of the training data.
@@
 Finally, the resulting merged model, with parameters denoted as $\theta_{\text{GRAPAM}}$, is used for inference on the entire children’s test set.
-This unified model is expected to more effectively capture speaker-group-specific characteristics, thereby improving recognition performance.
-IV Experimental setup
-IV-A Corpus
-TABLE I: My Science Tutor Children Speech Corpus statistics
+4 Experimental setup
+4.1 Corpus
+Table 1: My Science Tutor Children Speech Corpus statistics
 Training Validation Test
@@
 # of hours 117 18 19
-For our experiments, we used the My Science Tutor Corpus (MyST) dataset.
-MyST is one of the largest publicly accessible collections of English children’s speech, comprising approximately 400 hours.
-It encompasses dialogues between children and a virtual tutor across eight scientific domains, involving 1,372 students in grades three to five.
-The corpus is pre-partitioned, ensuring equitable representation of scientific domains and unique student occurrences within each partition.
-However, only 45% of utterances are transcribed at the word level.
-In this work, we excluded utterances shorter than one second, mainly containing silence, and those longer than 20 seconds, due to GPU constraints.
-After this filtering process, the dataset comprises 56,859 utterances from 738 speakers, totaling approximately 154 hours of speech.
-More detailed statistics of the different partitions are provided in Table I.
-IV-B Implementation details
-All experiments were conducted using the SpeechBrain toolkit [ravanelli2021speechbrain].
-For the ASR model, we employed Whisper-medium [^1] https://huggingface.co/openai/whisper-medium.en [radford2023robust], which has previously demonstrated strong performance when fine-tuned on children’s speech [jain23_interspeech].
-Whisper-medium is a Transformer-based encoder-decoder architecture comprising approximately 763.9M parameters, with 24 encoder layers and 24 decoder layers.
-The model was originally trained on a large-scale, multilingual dataset consisting of approximately 680k hours of transcribed speech.
-In this work, we evaluate three types of utterance-level representations for clustering.
-First, we extract 192-dimensional speaker embeddings using a pre-trained ECAPA-TDNN model [^2] https://huggingface.co/speechbrain/spkrec-ecapa-voxceleb.
-Second, we compute a set of 100-dimensional low-level speech metrics using the Librosa library [mcfee2015librosa], including pitch, pause rate, speaking rate, spectral centroid, and signal-to-noise ratio; these features are subsequently reduced to 16 dimensions using Principal Component Analysis (PCA).
-Third, we introduce a one dimensional WER-based representation derived from zero-shot inference WERs using the Whisper model, clustering utterances based on their predicted transcription quality.
-For all three representation types, we apply the $k$-means clustering algorithm from scikit-learn [pedregosa2011scikit], with a fixed number of clusters.
-We observed that, despite employing k-means, the cluster sizes remain balanced across all methods.
-Unless stated otherwise, we use $K=3$ clusters throughout our experiments.
-As for comparison, we also include random clustering, which assigns utterances to clusters uniformly at random.
-For Partial Fine-Tuning (PFT), we experiment with three configurations: fine-tuning (i) all model parameters (763.9M), (ii) only the attention components (302.4M), and (iii) only the feed-forward network (402.9M) (FFN) components across all Transformer layers.
-All experiments are conducted on a single NVIDIA RTX A6000 GPU (48 GB), with a batch size of 16, a learning rate of $1\times 10^{-5}$, and one epoch of training.
-The models are optimised using the negative log-likelihood (NLL) loss function.
-Finally, for model merging, we used LERP with uniform interpolation of the different models present in the configuration, ensuring that each group of children contributed equally to the final merged model.
-V Results and discussion
-To comprehensively assess the effectiveness of GRAPAM, we conduct a series of ablation studies aimed at disentangling the contributions of its core components.
-In particular, we evaluate the impact of different utterance clustering strategies, the effect of varying the number of groups, the number of selected merged models, the role of PFT, and alternative merging strategies including iterative model merging.
-These experiments are presented and analysed in detail in the remainder of this section.
-It is important to note that this study does not explore the influence of model size, using self-supervised learning (SSL) models instead of Whisper, or the volume of training data.
-These factors have already been examined in prior work on model merging for children’s ASR [shankar2025selective].
-V-A Group-aware model merging
-TABLE II: Results of different combinations for GRAPAM evaluated in WER (%).
-Values highlighted in grey indicate improvement relative to full-model fine-tuning on the entire dataset.
-The best-performing combinations are shown in bold.
-$\mathcal{M}$ Clustering method
-$\theta_{all}$ $\theta_{1}$ $\theta_{2}$ $\theta_{3}$ Spk-emb Random ZS WER LSM
-14.05
-x 9.95
-x 10.34 10.11 10.25 9.73
-x 9.79 9.69 10.32 10.61
-x 10.36 10.63 10.18 10.03
-x x 9.48 9.91 10.52 10.12
-x x 9.77 10.34 10.09 9.45
-x x 9.41 9.74 9.88 9.84
-x x x 9.68 9.68 10.01 9.41
-x x 9.72 9.98 10.32 9.79
-x x 9.41 9.67 9.86 9.83
-x x 9.76 9.77 9.96 9.70
-x x x 9.36 9.64 9.98 9.97
-x x x 9.68 9.98 9.72 9.68
-x x x x 9.65 9.65 9.93 9.63
-Average 9.59 9.84 10.03 9.74
-To evaluate the effectiveness of our proposed approach, we first examine whether clustering the training dataset followed by model merging improves performance in children’s ASR, and which utterance embedding strategy yield the best results.
-As shown in Table II, the baseline system achieves a Word Error Rate (WER) of 14.05%.
-Fine-tuning the full model on the entire dataset significantly reduces the WER to 9.95%.
-This result confirms that SFT effectively mitigates the adult-speech bias inherent in the original pre-trained Whisper model.
-Next, we partition $\mathcal{D}$ by group using several clustering strategies and apply model merging.
-For each configuration of $\mathcal{M}$, the different selected fine-tuned models used for merging are listed in the leftmost columns of Table II.
-Among all evaluated methods, grouping using speaker-embedding (Spk-emb) achieved the overall strongest performance, with an average WER of 9.59% and the lowest WER configuration with a score of 9.36%.
-The Low-level Speech Metrics (LSM) embeddings clustering also performs well but with less consistent results, reaching an average WER of 9.74% and a best configuration at 9.41%.
-Random clustering achieves 9.84% on average and 9.41% at best.
-Meanwhile, the Zero-Shot (ZS) WER clustering performs the least effectively with an average of 10.03% WER and best at 9.86%, showing no improvement over the baseline.
-In the configurations where all the fine-tuned models trained on all the different data clusters are merged together (final row of Table II), or $\mathcal{M}=\{\theta_{all},\theta_{1},\theta_{2},\theta_{3}\}$, Spk-emb and Random embedding yield slightly higher WERs (9.65%) compared with LSM (9.63%).
-Nevertheless, all of them outperform the fine-tuning baseline, indicating that model merging remains an effective strategy.
-Overall, these findings demonstrate that model merging enhances system performance relative to fine-tuning on the entire dataset and in particular, group-aware with speaker-embedding achieves the strongest and most consistent improvements.
-V-B Influence of the number of groups
-TABLE III: Influence of the number of clusters combination of all fine-tuned model ($\mathcal{M}=\{\theta_{all},\theta_{1},\dots,\theta_{k}\}$) and the best combination of $\mathcal{M}$ and results in WER (%).
+For our experiments, we use the My Science Tutor Corpus (MyST) dataset, one of the largest publicly available collections of English children’s speech ($\sim$400 hours).
+MyST contains child–virtual tutor dialogues spanning eight science domains and includes recordings from 1,372 students in grades 3–5.
+The dataset is released with predefined partitions that balance scientific domains and ensure that each student appears in only one split.
+Only 45% of utterances have word-level transcripts.
+We discard utterances shorter than 1 s, dominated by silence, and those longer than 20 s due to GPU constraints.
+After filtering, the resulting dataset contains 56,859 utterances from 738 speakers, totaling $\sim$154 hours of speech.
+Detailed partition statistics are reported in Table 1.
+4.2 Implementation details
+All experiments were performed with SpeechBrain [ravanelli2021speechbrain] using Whisper-medium.en [radford2023robust], a 24-layer encoder–decoder Transformer with 763.9M parameters, pre-trained on $\sim$680k hours of multilingual speech and shown effective for children’s ASR [jain23_interspeech].
+We cluster utterances using three representation types: (i) 192-d speaker embeddings from a pre-trained ECAPA-TDNN model [^1] https://huggingface.co/speechbrain/spkrec-ecapa-voxceleb, (ii) 16-d acoustic descriptors obtained by extracting 100 Librosa low-level metrics [mcfee2015librosa] and applying PCA, and (iii) a 1-d WER-based score from Whisper zero-shot predictions.
+For each representation, we apply k-means from scikit-learn [pedregosa2011scikit] (default K=3); we also report a random clustering baseline.
+For PFT, we compare updating (i) all parameters, (ii) attention modules only, or (iii) FFN modules only.
+Training uses a single RTX A6000 (48GB), batch size 16, learning rate $10^{-5}$, one epoch, and NLL loss.
+For merging, we use LERP with uniform weights so each cluster contributes equally to the merged model.
+5 Results
+Table 2: WER (%) results for GRAPAM.
+Left: clustering-method variants.
+Right: partial fine-tuning variants (SFT/PFT).
+Grey indicates improvement relative to full-model fine-tuning on the entire dataset; diagonal pattern indicates improvement over full SFT but not over the corresponding PFT on $\mathcal{D}$.
+Best values are in bold.
+$\mathcal{M}$ Clustering method Fine-tuning method
+$\theta_{all}$ $\theta_{1}$ $\theta_{2}$ $\theta_{3}$ Spk-emb Random ZS WER LSM Full SFT PFT FFN PFT ATTN
+14.05 014.05
+x 09.95 9.95 9.48 9.46
+x 10.34 10.11 10.25 9.73 10.34 10.03 10.00
+x 9.79 9.69 10.32 10.61 9.79 9.68 9.68
+x 10.36 10.63 10.18 10.03 10.36 9.58 9.71
+x x 9.48 9.91 10.52 10.12 9.48 9.52 9.61
+x x 9.77 10.34 10.09 9.45 9.77 9.69 9.68
+x x 9.41 9.74 9.88 9.84 9.41 9.44 9.56
+x x x 9.68 9.68 10.01 9.41 9.68 9.45 9.60
+x x 9.72 9.98 10.32 9.79 9.72 9.66 9.66
+x x 9.41 9.67 9.86 9.83 9.41 9.32 9.62
+x x 9.76 9.77 9.96 9.70 9.76 9.66 9.66
+x x x 9.36 9.64 9.98 9.97 9.36 9.33 9.68
+x x x 9.68 9.98 9.72 9.68 9.68 9.32 9.66
+PFT x x x - - - - - 9.36 9.51
+x x x x 9.65 9.65 9.93 9.63 9.65 9.31 9.38
+Average 9.59 9.84 10.03 9.74 9.59 9.47 9.60
+5.1 Group-aware model merging
+To assess our approach, we first test whether clustering $\mathcal{D}$ followed by model merging improves children’s ASR, and which utterance representation is most effective.
+As shown in Table 2, the baseline yields 14.05% WER, while full-model fine-tuning on $\mathcal{D}$ reduces WER to 9.95%, confirming that SFT mitigates the adult-speech bias of the pretrained Whisper model.
+We then cluster $\mathcal{D}$ using several embedding strategies and merge the resulting domain models according to $\mathcal{M}$.
+Speaker-embedding (Spk-emb) clustering performs best overall, achieving 9.59% average WER and a best configuration of 9.36%.
+Low-level speech metrics (LSM) is competitive but less consistent (9.74% average; 9.41% best), while Random clustering is weaker (9.84% average; 9.41% best).
+Zero-shot WER clustering performs worst (10.03% average; 9.86% best), failing to improve over the fine-tuning baseline.
+When merging all cluster-specific models ($\mathcal{M}=\{\theta_{all},\theta_{1},\theta_{2},\theta_{3}\}$), Spk-emb and Random reach 9.65% WER and LSM 9.63%; all remain better than the 9.95% fine-tuning baseline, supporting model merging as an effective consolidation strategy.
+5.2 Partial model merging
+In Table 2 we compare three fine-tuning strategies: full-model fine-tuning (Full SFT), partial fine-tuning of FFN blocks (PFT FFN), and partial fine-tuning of attention blocks (PFT ATTN).
+Fine-tuning on the full dataset $\mathcal{D}$ yields 9.95% WER with Full SFT, while partial fine-tuning performs better at 9.48% (FFN) and 9.46% (ATTN), despite fewer trainable parameters.
+Group-aware model merging further reduces WER across configurations, reaching best scores of 9.31% (PFT FFN) and 9.38% (PFT ATTN).
+On average, PFT FFN is strongest (9.47%), outperforming Full SFT (9.59%) and PFT ATTN (9.60%).
+When merging models trained on the full dataset, fully fine-tuned checkpoints slightly outperform partially fine-tuned ones (9.31% vs.
+9.36% for PFT–FFN and 9.51% for PFT–ATTN).
+The best PFT results arise when merging all cluster models, $\mathcal{M}=\{\theta_{\text{all}},\theta_{1},\theta_{2},\theta_{3}\}$.
+5.3 Influence of the number of groups
+Table 3: Influence of the number of clusters combination of all fine-tuned model ($\mathcal{M}=\{\theta_{all},\theta_{1},\dots,\theta_{k}\}$) and the best combination of $\mathcal{M}$ and results in WER (%).
 Number of clusters Combination of all Best combination
@@
 4 9.64 9.33
-We next evaluate model merging with varying numbers of clusters.
-Table III presents both the best-performing configurations of $\mathcal{M}$ and the results obtained by merging all fine-tuned group-based models ($\mathcal{M}=\{\theta_{all},\theta_{1},\theta_{2},\theta_{3}\}$) using the Spk-emb clustering method.
-The single-cluster setting corresponds to training on the entire dataset.
-For the best configurations, performance improves consistently as the number of clusters increases, with the lowest WER achieved using four clusters with a score of 9.33%.
-When all group-based models are merged, the combined system also delivers strong results, remaining competitive with the best configuration performance, achieving 9.42% with two clusters.
-These findings suggest that increasing cluster granularity allows the model to more effectively capture speaker-specific acoustic characteristics, leading to improved ASR performance when selected the right models to merge.
-V-C Partial model merging
-TABLE IV: WER (%) results of GRAPAM with partial fine-tuning.
-Grey indicates performance better than both full and partial fine-tuning on $\mathcal{D}$; Diagonal lines pattern indicates improvement over full fine-tuning but not over partial fine-tuning; The best-performing combination is shown in bold.
-$\mathcal{M}$ Fine-tuning method
-$\theta_{all}$ $\theta_{1}$ $\theta_{2}$ $\theta_{3}$ Full SFT PFT FFN PFT ATTN
-x 9.95
-Partial 9.48 9.46
-x 10.34 10.03 10.00
-x 9.79 9.68 9.68
-x 10.36 9.58 9.71
-x x 9.48 9.52 9.61
-x x 9.77 9.69 9.68
-x x 9.41 9.44 9.56
-x x x 9.68 9.45 9.60
-x x 9.72 9.66 9.66
-x x 9.41 9.32 9.62
-x x 9.76 9.66 9.66
-x x x 9.36 9.33 9.68
-x x x 9.68 9.32 9.66
-x x x x 9.65 9.31 9.38
-Partial x x x 9.36 9.51
-Average 9.59 9.47 9.60
-Table IV reports the WER performance of GRAPAM under three fine-tuning strategies: full model fine-tuning (Full SFT), partial fine-tuning of feed-forward modules (PFT FFN), and partial fine-tuning of attention modules (PFT ATTN).
-When applied to the entire dataset $\mathcal{D}$, full fine-tuning yields a WER of 9.95%.
-In contrast, partial fine-tuning achieves superior performance, with WERs of 9.48% (FFN) and 9.46% (ATTN), despite involving fewer trainable parameters.
-Applying group-aware model merging further improves performance across all configurations, achieving the lowest WERs of 9.31% with PFT FFN and 9.38% with PFT ATTN.
-On average, PFT FFN delivers the strongest results (9.47%), outperforming both Full SFT (9.59%) and PFT ATTN (9.60%).
-Furthermore, we examine whether merging fully fine-tuned models or partially fine-tuned models, both trained on the entire data set, yields greater benefits within the GRAPAM framework.
-As shown in the final two rows of Table IV, full fine-tuning yields slightly better performance, achieving a WER of 9.31%, compared to 9.36% for PFT–FFN and 9.51% for PFT–ATTN.
-Interestingly, the best-performing PFT configurations correspond to merging all fine-tuned models, i.e., $\mathcal{M}=\{\theta_{\text{all}},\theta_{1},\theta_{2},\theta_{3}\}$.
-V-D Integration of Clustering and Fine-Tuning Strategies
-In this section, we explore the integration of the previously discussed clustering methods and fine-tuning strategies, assessing whether performance can be enhanced by merging models across different strategies or by combining them sequentially across training stages.
-V-D1 Heterogeneous merging
-Table V presents the WER results of GRAPAM for all combination of PFT and utterance embedding when merging all models from all clusters, where $\mathcal{M}=\{\theta_{all},\theta_{1},\theta_{2},\theta_{3}\}$.
-The table also includes the heterogeneous merging scenarios (MERGE ALL line and column), where models fine-tuned using different embedding types and fine-tuning strategies are combined together.
-The results indicate that PFT–FFN consistently yields the strongest performance with a score of 9.36% WER.
-Merging over utterance embedding strategy also performs competitively, recording a WER of 9.38% for both speaker embeddings and 9.40% for LSM and 9.39% for random clustering.
-Importantly, merging across the different utterance embeddings or fine-tuning strategies does not significantly improves or degrades performance, suggesting that the merging process effectively preserves model capabilities.
-Notably, the MERGE ALL configuration, which merges all models across clustering types and fine-tuning strategies, achieves a WER of 9.32%, matching the best individual configurations and demonstrating the robustness of the GRAPAM framework under heterogeneous integration.
-TABLE V: WER (%) results of different combination of PFT and utterance embedding as well as the heterogeneous merging of them. The best-performing combination is shown in bold.
-Full SFT PFT FFN PFT Attn MERGE ALL
-Spk embedding 9.65 9.31 9.38 9.38
+Table 3 analyses the effect of the number of clusters (Spk-emb) on GRAPAM, reporting both (i) merging all group models $\mathcal{M}=\{\theta_{all},\theta_{1},\dots,\theta_{K}\}$ and (ii) the best-performing subset of $\mathcal{M}$.
+The K=1 setting (no clustering) matches full-data fine-tuning at 9.95% WER.
+For the best subsets, increasing K improves performance, reaching the lowest WER with four clusters (9.33%).
+Merging all group models remains competitive, achieving its best result with two clusters (9.42%).
+Overall, finer cluster granularity appears to better capture speaker-related variability, yielding gains when the merged model set is appropriately selected.
+5.4 Heterogeneous merging
+Table 4 reports GRAPAM results when merging all cluster models, $\mathcal{M}=\{\theta_{all},\theta_{1},\theta_{2},\theta_{3}\}$, across PFT variants and utterance-embedding choices.
+It also includes heterogeneous settings (ALL row/column), where models trained with different clustering representations and fine-tuning strategies are merged.
+Overall, PFT–FFN yields the lowest WER (9.36%).
+Merging across embedding types remains competitive, with WERs of 9.38% (speaker embeddings), 9.40% (LSM), and 9.39% (random).
+Importantly, combining models across embeddings or fine-tuning strategies produces only minor changes, indicating that merging largely preserves the underlying capabilities of the constituent models.
+The full heterogeneous merge (ALL) achieves 9.32% WER, matching the best individual configurations and demonstrating robustness to heterogeneous merge.
+Table 4: WER (%) results of different combination of PFT and utterance embedding as well as the heterogeneous merging of them. The best-performing combination is shown in bold.
+Full SFT PFT FFN PFT Attn ALL
+Spk-emb 9.65 9.31 9.38 9.38
 LSM 9.63 9.32 9.45 9.40
 Random 9.65 9.64 9.38 9.39
-MERGE ALL 9.60 9.36 9.51 9.32
-V-D2 Iterative Group-Aware Partial Merging
-TABLE VI: WER (%) for different embedding methods (Spk embedding, LSM, Random) across PFT (FFN, ATTN, FULL) and iterations. Selected models as a source for the next turn are presented in bold.
-PFT Utterance embedding TURN 1 TURN 2 TURN 3
-Baseline - 9.48 10.14 10.28
+ALL 9.60 9.36 9.51 9.32
+5.5 Iterative Group-Aware Partial Merging
+Table 5: WER (%) for different embedding methods (Spk embedding, LSM, Random) across PFT (FFN, ATTN, FULL) and iterations. Selected models as a source for the next turn are presented in bold.
+PFT Utt embedding Iter 1 Iter 2 Iter 3
+Partial - 9.48 10.14 10.28
 FFN Spk embedding 9.31 9.30 9.35
@@
 Random 9.65 9.88 9.92
-Table VI presents the WER results obtained from successive iterations of GRAPAM, where, at each step, the best-performing configuration from the previous iteration is used as the pre-trained model for subsequent merging.
-In this experiment we used $\mathcal{M}=\{\theta_{all},\theta_{1},\theta_{2},\theta_{3}\}$.
-First, we observe that when the full model is trained for additional epochs on the entire dataset, signs of overfitting emerge, as indicated by an increase in WER from 9.48% to 10.28%.
-In contrast, the Iterative GRAPAM results demonstrate that initialising with speaker-embedding clustering and partial fine-tuning of the FFN yields the lowest WER of 9.28%, achieved in iteration 2 using LSM-based clustering and full fine-tuning.
-We hypothesise that this improvement arises from the model’s exposure to heterogeneous information across different cluster types at each iteration, enabling the progressive development of richer and more generalisable representations.
-However, beyond iteration 2, no further improvements are observed.
-This plateau suggests that the model has saturated the informative capacity of the available cluster partitions and may be beginning to overfit.
-These findings confirm that iterative group-aware partial model merging effectively enhances ASR performance.
-VI Conclusion and future work
-In this work, we presented GRAPAM, a parameter-efficient framework for adapting adult-pretrained ASR models to children’s speech through group-aware partial model merging.
-By combining unsupervised clustering, partial fine-tuning, and parameter-level interpolation, GRAPAM achieves a WER reduction from 9.95% to 9.31% on the MyST corpus.
-This new fine-tuning strategy for children’s ASR achieves comparable or better performance without needing multi-task training or additional data.
-Future work can extend GRAPAM in several directions.
-First, adaptive weighting, where interpolation coefficients are optimised on held-out validation data rather than set uniformly is a promising direction.
-Second, more advanced clustering techniques can be considered, potentially incorporating self-supervised speech representations to better capture age and speaker-related variability.
-Finally, GRAPAM could be extended to other domains like pathological or multilingual speech recognition.
+Table 5 reports WER for successive GRAPAM iterations, where the best configuration from one iteration is used to initialise the next, with $\mathcal{M}=\{\theta_{all},\theta_{1},\theta_{2},\theta_{3}\}$.
+Training the full model for additional epochs on the entire dataset degrades performance, with WER increasing from 9.48% to 10.28%, consistent with overfitting.
+In contrast, iterative GRAPAM improves recognition: starting from speaker-embedding clustering with FFN partial fine-tuning, the best result is 9.28%, obtained at iteration 2 after applying LSM-based clustering with full fine-tuning.
+This gain is consistent with incremental exposure to complementary cluster structures across iterations, which may encourage more robust representations.
+After iteration 2, performance saturates and no further gains are observed, suggesting limited additional information in the available partitions and potential onset of overfitting.
+Overall, the results support iterative group-aware model merging as an effective mechanism for improving children’s ASR.
+6 Conclusion and future work
+In this work, we introduced GRAPAM, a parameter-efficient framework for adapting adult-pretrained ASR models to children’s speech via group-aware partial model merging.
+By integrating unsupervised clustering, partial fine-tuning, and parameter-level interpolation, GRAPAM reduces WER on MyST from 9.95% to 9.31%, achieving comparable or improved accuracy without multi-task training or additional data.
+Several directions remain for future work.
+These include replacing uniform interpolation with adaptive weighting, learning merge coefficients on held-out validation data.
+Exploring stronger merging strategies, including clustering based on self-supervised speech representations and extending GRAPAM to other tasks such as pathological speech.
+7 Acknowledgements
+This work has been submitted to Interspeech 2026.
+Work supported by Portuguese national funds through Fundação para a Ciência e a Tecnologia (FCT), with references UIDB/50021/2020 and 2022/12328/BD, as well as by the Portuguese Recovery and Resilience Plan (RRP) through project C644865762-00000008 (Accelerat.AI).